# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 4.1.5. Feature Relevancy
Feature relevance refers to identifying and selecting the most informative and significant features or variables that contribute to the model’s predictive power. In a dataset, various features or variables are typically collected for each instance, representing different aspects or characteristics of the data. However, not all features may be equally relevant or valuable for the AI model. Feature relevance aims to identify the subset of features that are most influential in making accurate predictions or capturing the underlying patterns in the data.
@@
Example: Consider a dataset for predicting house prices. The dataset may include the number of bedrooms, square footage, location, and proximity to amenities. In this case, feature relevance or selection would involve analyzing the relationship between each feature and the target variable (house prices) to determine which ones have the strongest correlation or impact on the predictions. Features that are found to have weak or negligible influence can be excluded, resulting in a more concise and informative feature set.
@@
Metrics in Literature: The column heterogeneity measure proposed by Dai et al. (Dai et al., 2006) uses soft clustering techniques and mutual information to quantify the relevancy of features in a dataset. Soft clustering assigns fractional memberships to data points across multiple clusters, allowing for a more nuanced representation of underlying patterns. Mutual information is used to measure the dependence between feature values and soft clustering results, capturing the association between them. The computed mutual information values are then used to derive column heterogeneity scores for each feature, indicating relevancy and diversity. Features with higher scores are considered more informative and likely to contribute significantly to AI tasks.
@@
The survey by Li et al. (Li et al., 2017) collectively explores similarity-based feature selection metrics including Laplacian Score, SPEC, Fisher Score, Trace Ratio, and ReliefF. The Laplacian Score algorithm by He et al. (He et al., 2005) constructs affinity and Laplacian matrices to gauge similarities and differences among data points, producing scores that prioritize features capturing underlying data structures. SPEC, an extension of Laplacian Score, introduced by Zhao et al. (Zhao and Liu, 2007), emphasizes alignment with data structure through spectral analysis, adapting to unsupervised and supervised scenarios. On the other hand, Duda et al’s (Duda et al., 2012) Fisher Score emphasizes comparability within classes and distinctiveness between classes, while the Trace Ratio criterion by Nie et al.(Nie et al., 2008) and ReliefF algorithm by Robnik-Sikonja et al. (Robnik-Šikonja and Kononenko, 2003) emphasize within-class similarity and between-class dissimilarity. Collectively, these algorithms highlight the significance of exploiting data relationships and class structures.
Li et al. (Li et al., 2017) focus on information-theoretical-based feature selection methods, including Mutual Information Maximization (MIM) (Lewis (Lewis, 1992)). MIM relies on the concept of entropy to evaluate the significance of features by measuring the reduction in uncertainty they bring to a classification task. MIM evaluates each feature’s significance based on its correlation with class labels, aiming to identify features strongly associated with the target variable. Features with higher Mutual Information (MI) scores are considered more informative and are selected until the desired number of features is reached.
@@
Li et al. (Li et al., 2017) comprehensively survey prominent statistical-based methods, outlining their roles and applications. Among these techniques, the Low Variance method measures feature relevance by evaluating variances and removing features with variances below a specified threshold. In binary classification, the T-Score method, proposed by Davis et al. (Davis and Sampson, 1986), quantifies a feature’s capacity to differentiate classes by calculating T-scores based on class means and standard deviations, with higher scores indicating stronger discriminatory power. Conversely, the Chi-Square Score method, introduced by Liu et al. (Liu and Setiono, 1995), assesses feature-class independence through an independence test derived from differences between observed and expected frequencies. Gini Index (Gini, 1912) evaluates a feature’s partitioning potential across different classes, using class probabilities, considering how effectively its values divide the dataset. Another method, CFS (Correlation-based Feature Selection) by Hall et al. (Hall and Smith, 1999), evaluates feature subsets worth using a correlation-based heuristic. The CFS score balances predictive power with redundancy, using symmetrical uncertainty to assess feature correlations.
Gupta et al. (Gupta et al., 2021) have showcased the practical application of the feature relevance metric as a valuable data readiness measure in their data quality toolkit. This metric effectively evaluates the importance of each feature and assigns a relevance score to rank them accordingly. The quality metric component of this measure plays a pivotal role in identifying and suggesting the elimination of less relevant features, ensuring that only the most pertinent ones are retained for subsequent classification tasks.
Summary: Feature relevancy helps identify and select the most informative and significant features that contribute to the predictive power of AI models. Dai et al. introduce a column heterogeneity measure using soft clustering and mutual information. Li et al.’s survey explores similarity-based, information-theoretical, and statistical methods. These include Laplacian Score, SPEC, Fisher Score, Trace Ratio, ReliefF, Mutual Information Maximization (MIM), Low Variance, T-Score, Chi-Square Score, Gini Index, and CFS. Gupta et al.’s toolkit offers a feature relevance metric that assigns scores, aiding in feature selection for optimized model performance.
@@ 4.1.6. Class Imbalance
Class imbalance in the context of AI-ready data refers to a situation where the distribution of instances among different classes in a dataset is significantly skewed or uneven. It means that one or more classes have a much larger number of instances than others, leading to an imbalanced representation of the classes.
Example: Consider a dataset for customer churn prediction in a telecommunications company. The dataset contains information about customer attributes such as age, gender, usage patterns, and whether or not the customer churned. In this case, the positive class refers to the smaller group or category within the dataset, consisting of customers who churned, while the negative class denotes the larger group or category in the dataset, consisting of customers who did not churn. If only 5% of the instances belong to the positive churn class, while the remaining 95% belong to the negative non-churn class, it indicates a severe class imbalance.
@@
Class imbalance can pose challenges during AI model training and evaluation. Models trained on imbalanced data tend to prioritize the majority class, resulting in poor performance for the minority class. In the customer churn example, an imbalanced dataset may lead to a model that performs well in predicting non-churn instances but performs poorly in identifying churn instances, which are often the class of interest.
@@
Metrics in Literature: The Individual Bayes Imbalance Impact Index (IBI3), introduced by Lu et al. (Lu et al., 2019) assesses the impact of class imbalance on individual samples, providing insights into potential biases and dataset limitations. IBI3 quantifies the difference in posterior probabilities between balanced and imbalanced scenarios, revealing how class imbalance influences classification outcomes. It requires trained models and estimation of posterior probabilities to calculate, making it essential to have access to both for an accurate assessment. IBI3 measures the influence of class imbalance on classification outcomes for each minority class sample, with lower values indicating less impact and higher values indicating a more significant impact.
@@
The Imbalance Ratio (IR), introduced by Francisco et al. (Alberto et al., 2018), is a widely used metric to quantify the level of class imbalance in a dataset, especially in binary classification problems. It provides a numerical representation of the discrepancy between the majority class instances and minority class instances in the dataset. The IR is calculated by dividing the count of instances in the majority class ($N\textunderscore majority$) by the count of instances in the minority class ($N\textunderscore minority$). A higher IR indicates a more imbalanced dataset with a significant disparity between class frequencies, while a lower IR suggests a relatively balanced or less imbalanced dataset.
@@
Ortigosa-Hernández et al. (Ortigosa-Hernández et al., 2017) propose the Imbalance Degree (ID) as an advanced metric to measure class imbalance, considering specific characteristics of the class distribution for a more accurate representation. Despite its advantages, ID has drawbacks, such as sensitivity to the choice of distance function and potential unreliability in extreme cases. In contrast, Zhu et al. (Zhu et al., 2018) introduce the Likelihood Ratio Imbalance Degree (LRID) as a novel metric that overcomes ID limitations. LRID uses the likelihood ratio (LR) test, providing a high-resolution measurement of imbalance by comparing empirical class distribution to a balanced distribution. LRID offers insights into the deviation from balance, aiding researchers in making informed decisions on data preprocessing to mitigate imbalance’s impact on AI model performance.
In addition to LRID, Gupta et al. (Gupta et al., 2021) propose the class parity metric, which considers various data properties, including the imbalance ratio, dataset size, and proportion of difficult samples in the extreme minority class. By analyzing these factors, data scientists can gain a comprehensive understanding of class imbalance in the dataset and apply remedial re-sampling strategies to balance the data effectively.
@@
Summary: Class imbalance in AI-ready data is assessed using metrics like Imbalance Ratio (IR), with Imbalance Degree (ID) offering nuanced measurements. Likelihood Ratio Imbalance Degree (LRID) provides high-resolution assessment through the likelihood ratio test. Gupta et al.’s class parity metric considers imbalance ratio, dataset size, and difficult samples for comprehensive evaluation and mitigation.
@@ 4.1.7. Class Separability
In the context of AI, it refers to the degree of similarity or shared characteristics between different classes or categories within the dataset. It indicates how much the data points from different classes overlap or share common features, making it challenging for an AI system to distinguish between them accurately.
Example: Consider a facial recognition dataset consisting of two classes: “smiling” and “not smiling.” The overlap in this dataset refers to the extent to which the facial features of individuals in the “smiling” class overlap with those in the “not smiling” class. If the dataset contains many instances where individuals in both classes have similar facial expressions or features, it indicates a high overlap. In this case, it becomes more difficult for the AI system to correctly classify whether a person is smiling or not based solely on facial features.
Metrics in Literature: Gupta et al.’s data quality toolkit introduces a class overlap metric, which identifies and quantifies overlapping regions among different classes in a dataset, analyzing data points residing in overlapping regions of the data space. Additionally, the evaluation of class overlap in imbalanced classification settings is addressed through metrics such as the $R$-value and augmented $R$-value. Sejong’s (Oh, 2011) $R$-value assesses the extent of overlap between classes by considering the proportion of instances in a specific class located in regions of the feature space shared with instances from other classes. Borsos et al. Borsos et al. (Borsos et al., 2018) enhance this approach with the augmented $R$-value, which incorporates the dataset’s imbalance ratio (IR), providing a weighted measure that combines class overlap and dataset imbalance for a more comprehensive understanding.
Summary: Class separability pertains to the level of similarity among diverse classes in a dataset. Gupta et al.’s data quality toolkit introduces a metric to detect overlapping areas between classes, assessing data points that are close yet belong to different classes. The R-value is a measure proposed to quantify the degree of overlap in imbalanced classification problems. By considering the R-value of each class and the dataset’s imbalance ratio, an augmented R-value can be calculated, providing a more comprehensive measure of overlap.
@@ 4.1.8. Discrimination Index
Mitigating discrimination in data refers to the process of understanding and addressing potential biases or discriminatory outcomes that may arise when using data in AI systems. It involves recognizing and mitigating unfairness or unjust treatment towards certain individuals or groups that may be encoded or perpetuated in the data used to train or deploy AI models.
@@
Example: An example to illustrate fairness awareness in data for AI is in the context of hiring practices. Consider a company that uses an AI system to screen job applicants based on their resumes. The AI model is trained on historical data of successful candidates and their qualifications. However, suppose the historical data is biased and reflects discriminatory practices, such as favoring candidates from certain gender or ethnic groups. In that case, the AI model may inadvertently perpetuate those biases and lead to unfair outcomes in the hiring process.
@@
Metrics in Literature: The Difference metric, introduced by Azzalini et al. (Azzalini et al., 2022), assesses the degree of unethical behavior or bias within a dependency by comparing the confidence of that dependency with and without consideration of sensitive attributes. A higher Difference value indicates a stronger indication of unfair behavior, with the support of the Approximate Conditional Functional Dependency (ACFD) also factored in. Additionally, the authors propose the P-Difference metric to address dependencies involving multiple sensitive attributes, measuring unfairness tied to specific attributes. The P-Difference gauges the impact on dependency confidence by excluding one sensitive attribute at a time, highlighting influential attributes contributing to unfairness.
@@
Feldman et al. (Feldman et al., 2015) introduce the “Likelihood Ratio” ($LR_{+}$) metric to measure disparate impact in a dataset, calculated based on sensitivity and specificity. It assesses the impact of the protected class on classification outcomes, but it requires a model trained on the dataset to generate results. Celis et al. (Celis et al., 2020) contribute two metrics for measuring discrimination based on sensitive attributes: the ”representation rate” quantifies fairness in representing different attribute values, assessing the ratio’s adherence to a threshold, while the ”statistical rate” evaluates fairness through conditional probabilities of class labels given attribute values, identifying potential discrimination. These metrics provide tailored quantitative fairness evaluation, offering flexibility based on specific application requirements.
@@
Simonetta et al. (Simonetta et al., 2021) introduce two distinct metrics that contribute to assessing fairness, bias, and completeness in AI-ready datasets. The first metric, a “combinatorial metric,” evaluates dataset completeness by focusing on the distinct combinations of categories within specific columns. It quantifies completeness by comparing the total count of unique data points to the expected number of distinct combinations. This metric provides insights into how well the dataset represents the full spectrum of category combinations. In contrast, the second metric, rooted in “frame theory,” offers a sophisticated approach to measuring fairness and bias. It treats the dataset as a matrix and applies operations to analyze the distribution of vectors within the matrix. Eigenvalues obtained from this matrix assessment gauge the tightness of the frame, with uniform eigenvalue distribution indicating a balanced dataset. The Gini-Simpson index ((Simpson, 1949)) is used to further assess balance and homogeneity. The difference lies in the focus: the combinatorial metric targets the representation of distinct combinations, while the frame theory-based metric centers on the overall distribution and balance of the dataset’s vectors.
The Amazon SageMaker Developer Guide (Kemka, 2019) introduces the demographic disparity metric (DD) to assess imbalanced proportions of accepted and rejected outcomes within demographic facets as a pre-training data bias measure. The conditional form (CDD) addresses Simpson’s paradox by analyzing subgroups, revealing apparent imbalances. In contrast, Gupta et al.’s data quality toolkit presents a disparate impact measure to quantify group discrimination, offering a score for assessing fairness. The toolkit includes remediation strategies to mitigate bias in data and its impact on AI model applications, enhancing group fairness.
@@
Summary: The discrimination index allows analysts to quantify and measure biases or discriminatory outcomes encoded in the data used for training and deploying AI models. The metrics such as the Difference metric, P-Difference metric, Likelihood Ratio ($LR_{+}$), representation rate, statistical rate, completeness metric, and frame theory-based metrics, provide quantitative measures to detect and evaluate unfairness, disparate impact, and discriminatory behavior in the dataset.

### v2

@@ 3.1.5. Feature Relevance
This refers to identifying and selecting the most informative features or variables that contribute to an AI model. In a dataset, various features or variables are typically collected for each instance, representing different aspects or characteristics of the data. However, not all features may be equally relevant or valuable for an AI model. Feature relevance metric aims to identify the subset of features that are most influential in making accurate predictions or capturing the underlying patterns in the data.
@@
Example: Consider a dataset for predicting housing prices that includes number of bedrooms, square footage, location, schools, and proximity to amenities. In this case, feature relevance would involve analyzing the relationship between each feature and the target variable (house prices) to determine which ones have the strongest correlation or impact on the predictions. Features that have weak or negligible influence can be excluded.
@@
Metrics in Literature: The column heterogeneity measure proposed by Dai et al. (2006) uses soft clustering techniques and mutual information to quantify the relevance of features in a dataset. Soft clustering assigns fractional memberships to data points across multiple clusters. Mutual information measures the dependence between feature values and soft clustering results, capturing their association. The computed mutual information values are then used to derive column heterogeneity scores for each feature. Features with higher scores are considered more informative.
@@
A survey by Li et al. (2017) collectively explores similarity-based feature selection metrics, including Laplacian Score, SPEC, Fisher Score, Trace Ratio, and ReliefF. The Laplacian Score algorithm by He et al. (2005) constructs affinity and Laplacian matrices to measure similarities and differences among data points, producing scores that prioritize features capturing underlying data structures. SPEC, an extension of the Laplacian Score, introduced by Zhao and Liu (2007), emphasizes alignment with data structure through spectral analysis. Duda et al.’s (Duda et al., 2012) Fisher Score emphasizes comparability within classes and distinctiveness between classes, while the Trace Ratio criterion by Nie et al. (2008) and ReliefF algorithm by Robnik-Šikonja and Kononenko (2003) emphasize within-class similarity and between-class dissimilarity. Collectively, these algorithms highlight the significance of exploiting data relationships and class structures.
@@
Li et al. (2017) focus on information-theory-based feature selection methods, including Mutual Information Maximization (MIM) (Lewis (1992)). MIM relies on the concept of entropy to evaluate the significance of features by measuring the reduction in uncertainty they bring to a classification task. MIM evaluates each feature’s significance based on its correlation with class labels. Features with higher Mutual Information (MI) scores are considered more informative and are selected until the desired number of features is reached. Li et al. (2017) also provides statistical methods, highlighting their roles and applications in various fields. Among these methods, the Low Variance method measures feature relevance by evaluating variances and removing features with variances below a specified threshold. In binary classification, the T-Score method, proposed by Davis and Sampson (1986), quantifies a feature’s capacity to differentiate classes by calculating T-scores based on class means and standard deviations, with higher scores indicating stronger discriminatory power. Conversely, the Chi-Square Score method, introduced by Liu and Setiono (1995), assesses feature-class independence through an independence test derived from differences between observed and expected frequencies. Gini Index (Gini, 1912) evaluates a feature’s partitioning potential across different classes, using class probabilities, considering how effectively its values divide the dataset. The Correlation-based Feature Selection (CFS) by Hall and Smith (1999) evaluates feature subsets worth using a correlation-based heuristic. The CFS score balances predictive power with redundancy using symmetrical uncertainty.
Impact on AI: By identifying the most important features, AI models can concentrate on key data aspects, which reduces noise and computational demands and enhances performance. This process helps mitigate the “curse of dimensionality,” speeds up training, and can prevent over-fitting. It also enhances model interpretability by highlighting key factors driving predictions or decisions.
Summary: Feature relevance helps in identifying and selecting the most informative and significant features that contribute to the predictive power of AI models. Existing feature relevance metrics use statistical techniques such as soft clustering, similarity, and information theory.
@@ 3.1.6. Class Imbalance
In the context of data for AI, class imbalance refers to the highly skewed or uneven distribution of instances among different classes (categories) in a dataset. It means that one or more classes appear more than others, leading to an imbalanced representation.
Example: Class imbalance often appears in datasets with rare event detection, such as credit fraud detection, earthquake prediction, network intrusion detection, customer churn prediction, rare disease diagnosis, rare species of animal sightings, etc. In these cases, rare events or anomalies are the focus of detection.
@@
Class imbalance can pose challenges during AI model training and evaluation. Models trained on imbalanced data tend to prioritize the majority class, resulting in poor prediction performance for the minority class. In an anomaly detection example, an imbalanced dataset may lead to a prediction model that performs well in predicting normal instances but performs poorly in identifying anomalies, which are often the class of interest.
@@
Metrics in Literature: The Individual Bayes Imbalance Impact Index (IBI3), introduced by Lu et al. (2019) assesses the impact of class imbalance on individual samples, providing insights into potential biases and dataset limitations. IBI3 quantifies the difference in posterior probabilities between balanced and imbalanced scenarios, revealing how class imbalance influences classification outcomes. It requires trained models and estimation of posterior probabilities to calculate, making it essential to have access to both for an accurate assessment. IBI3 measures the influence of class imbalance on classification outcomes for each minority class sample, with lower values indicating less impact.
@@
Imbalance Ratio (IR), introduced by Alberto et al. (2018), is a widely used metric to quantify the level of class imbalance in a dataset, especially in binary classification problems. It provides a numerical representation of the discrepancy between the majority and minority class instances. IR is calculated by dividing the count of instances in the majority class ($N\textunderscore majority$) by the count of instances in the minority class ($N\textunderscore minority$). A higher IR indicates a more imbalance.
@@
Ortigosa-Hernández et al. (2017) propose Imbalance Degree (ID) as a metric to measure class imbalance, considering specific characteristics of the class distribution. Despite its advantages, ID has drawbacks, such as sensitivity to the choice of distance function and potential unreliability in extreme cases. In contrast, Zhu et al. (2018) introduce the Likelihood Ratio Imbalance Degree (LRID) to overcome ID’s limitations. LRID uses the likelihood ratio (LR) test, providing a high-resolution measurement of imbalance by comparing empirical class distribution to a balanced distribution. Gupta et al. (et al., 2021) propose the class parity metric, which considers various data properties, including the imbalance ratio, dataset size, and proportion of difficult samples in the extreme minority class.
Impact on AI: Class imbalance in datasets can significantly impact AI systems, particularly in classification tasks. Class imbalance can cause models to overlook important patterns in minority classes, potentially leading to misclassification.
@@
Summary: Class imbalance in AI-ready data is assessed using metrics like Imbalance Ratio (IR), with Imbalance Degree (ID) offering nuanced measurements. Likelihood Ratio Imbalance Degree (LRID) provides a high-resolution assessment through the likelihood ratio test. Gupta et al.’s class parity metric considers imbalance ratio, dataset size, and difficult samples.
@@ 3.1.7. Class Separability
Class Separability refers to the degree of similarity or shared characteristics between different classes or categories within the dataset. It measures the overlap or sharing of common features among the data points from different classes.
Example: Consider a facial recognition dataset consisting of two classes: “smiling” and “not smiling.” The overlap in this dataset refers to the extent to which the facial features of individuals in these two classes. If the dataset contains many instances where individuals in both classes have similar facial expressions or features, it indicates a high overlap.
Metrics in Literature: Gupta et al.’s Data Quality Toolkit (DQT) (Developer, 2021) introduces a class overlap metric, which quantifies overlapping regions among different classes in a dataset by analyzing data points in overlapping regions of the data space. Additionally, the evaluation of class overlap in imbalanced classification settings is addressed through metrics such as the $R$-value and augmented $R$-value. Sejong’s (Oh, 2011) $R$-value assesses the extent of overlap between classes by considering the proportion of instances in a specific class located in regions of the feature space shared with instances from other classes. Borsos et al. (2018) enhance this approach with the augmented $R$-value, which considers the dataset’s imbalance ratio (IR). It provides a weighted measure that combines class overlap and dataset imbalance for a more comprehensive understanding.
Impact on AI: Class separability impacts AI systems in classification tasks by influencing a model’s ability to accurately distinguish between different categories. The model can establish clear decision boundaries when classes are well-separated in the feature space. This leads to improved accuracy, faster training, better generalization to new data, and increased resilience against noise. This also enhances model interpretability, making the decision-making process more transparent. Conversely, low-class separability can lead to more misclassifications.
Summary: Class separability is the level of similarity among diverse classes in a dataset. DQT introduces a metric to detect overlapping areas between classes, assessing data points that are close yet belong to different classes. The R-value is a measure proposed to quantify the degree of overlap in imbalanced classification problems.
@@ 3.1.8. Discrimination Index
Discrimination in data refers to biases that may cause discriminatory outcomes in AI systems. It measures unfairness or unjust treatment towards individuals or groups that may be encoded in the data.
@@
Example: Consider a company that uses an AI system to filter job applicants based on their resumes. The AI model is trained on historical data of successful candidates and their qualifications. If the historical data is biased and reflects discriminatory practices, such as favoring candidates from certain demographic groups, the AI model may unknowingly sustain those biases and lead to unfair outcomes in the hiring process.
@@
Metrics in Literature: The Difference metric, introduced by Azzalini et al. (2022), assesses the degree of bias within a dependency by comparing the confidence of that dependency with and without consideration of sensitive attributes. A higher Difference value indicates a stronger indication of unfair behavior. It is further supported by the Approximate Conditional Functional Dependency (ACFD). Additionally, the authors propose the P-Difference metric to measure the impact on dependency confidence by excluding one sensitive attribute at a time. This highlights the influential attributes contributing to unfairness.
@@
Feldman et al. (2015) introduce the “Likelihood Ratio” ($LR_{+}$) metric to measure disparate impact in a dataset, calculated based on sensitivity and specificity. It assesses the impact of the protected class on classification outcomes, but it requires a model trained on the dataset to generate results. Celis et al. (2020) introduce two metrics for assessing discrimination based on sensitive attributes. The “representation rate” measures fairness by checking how well different attribute values are represented compared to a set threshold. The “statistical rate” evaluates fairness by analyzing the conditional probabilities of class labels given attribute values, helping to identify potential discrimination. These metrics provide quantitative fairness evaluation, offering flexibility based on specific application requirements.
@@
Simonetta et al. (2021) introduce two metrics that contribute to assessing fairness, bias, and completeness. The first metric, a “combinatorial metric,” evaluates dataset completeness by focusing on the distinct combinations of categories within specific columns. It quantifies completeness by comparing the total count of unique data points to the expected number of distinct combinations. In contrast, the second metric, based on “frame theory,”(Daubechies, 1992) offers a sophisticated approach to measuring fairness and bias. It treats the dataset as a matrix and applies operations to analyze the distribution of vectors within the matrix. Eigenvalues obtained from this matrix assessment measure the tightness of the frame, with uniform eigenvalue distribution indicating a balanced dataset. The Gini-Simpson index ((Simpson, 1949)) is used to assess balance and homogeneity further. The combinatorial metric targets the representation of distinct combinations, while the frame theory-based metric considers the overall distribution and balance of the dataset’s vectors.
The Amazon SageMaker Developer Guide (Kemka, 2019) uses various metrics for identifying bias in data. Class Imbalance (CI) measures sample distribution across the sensitive attributes, and Difference in Proportions of Labels (DPL) assesses outcome disparities. The guide also leverages various divergence metrics like Kullback-Leibler (KL), Jensen-Shannon (JS), and Lp-norm evaluate differences in the outcome distributions across demographic facets. Total Variation Distance (TVD) and Kolmogorov-Smirnov (KS) measure the degree of distribution divergence, and Conditional Demographic Disparity (CDD) assesses outcome disparities within subgroups. In contrast, DQT (Developer, 2021) presents a disparate impact measure to quantify group discrimination, offering a score for assessing fairness. DQT also includes remediation strategies to mitigate bias in data.
Impact on AI: Unfair data can significantly impact AI systems by leading to biased and incorrect decisions. When AI is trained on biased or unrepresentative data, it can produce outcomes that systematically disadvantage certain groups based on race, gender, socioeconomic status, or other factors. This can result in discriminatory practices in critical areas such as hiring, lending, healthcare, and law enforcement.
@@
Summary: The discrimination index allows analysts to quantify and measure biases or discriminatory outcomes encoded in the data used for training and deploying AI models. The metrics, such as the Difference metric, P-Difference metric, Likelihood Ratio ($LR_{+}$), representation rate, statistical rate, completeness metric, divergence metrics, and frame theory-based metrics provide quantitative measures to detect discriminatory behavior in the dataset.